Circular No. 6164

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

POSSIBLE MICROLENSING IN THE GALACTIC BULGE A. Udalski and M. Szymanski, Warsaw University Observatory, on be-

half of the OGLE microlensing search team, communicate: “The OGLE ‘Early Warning System’ (cf. IAUC 5997; Udalski et al. 1994, Acta Astron. 44, 227) has detected an object in the galactic bulge that fulfills the criteria of a ‘prime microlensing candidate’. Observations collected since Apr. 11 (the beginning of the OGLE 1995 season) with the 1-m Swope telescope at the Las Campanas Observatory have shown that the star (α = 18^{h}06^{m}48^{s}.68, δ = −26^{o}37′24′′.0, equinox 2000.0) had brightened ∼ 0.6 mag, while in the 1993 and 1994 observing seasons the star was constant at I = 19.08 and V − I = +2.3. On Apr. 18.25 UT, its brightness was I ∼ 18.65. The star seems to be on the declining branch of the microlensing lightcurve, 3–5 days after maximum. However, due to the short period of observations in the 1995 season, we are not yet able to determine the microlensing lightcurve parameters, and we cannot rule out the possibility that the brightening is related to some other kind of variability. We strongly encourage follow-up observations in many wavelength pass-bands to confirm the microlensing hypothesis. A finding chart, lightcurve, and regularly updated photometry can be found in the OGLE archive (anonymous ftp

sirius.astrouw.edu.pl, directory /ogle/ews/1995/mm1-a).”

AL COMAE BERENICES D. Nogami and T. Kato, Department of Astronomy, Kyoto Univer-

sity, write: “V -band CCD photometry obtained with a 0.60-m reflector at Ouda Station, Kyoto University, indicates that the humplike feature reported independently by Patterson and by DeYoung (IAUC 6157) showed a continous decay until Apr. 12. The data between Apr. 17.440 and 17.691 UT clearly demonstrate a periodic modulation with a period of 0.0576 ± 0.0002 day and an amplitude of 0.18 mag. This period is slightly longer than those reported earlier and one (0.05666 ± 0.00002 day) determined by us from Apr. 7–12 observations. This fact indicates that the new features are genuine superhumps, and earlier detection may reflect a variation related to the orbital motion of the underlying binary. The present observation thus first firmly identifies the SU UMa-type nature of AL Com. However, the baseline is not so long. Follow-up observations are expected.”

1995 April 19 Daniel W. E. Green